OpenStudy (anonymous):

What is the least common denominator of the fraction 2 over x plus the fraction x over x plus 1 plus the fraction x plus 1 over 3 x

OpenStudy (anonymous):

\[\frac{ 2 }{ x } + \frac{ x }{ x+1 } + \frac{ x+1 }{ 3x }\]

OpenStudy (anonymous):

I'll give a medal.

OpenStudy (anonymous):

Can u help me rob?